# RealWebAssist: A Benchmark for Long-Horizon Web Assistance with Real-World Users

### v1

@@ 4.2 Results
| VLM + Grounding | GPT-4o | 10.3 | 21.7 | 66.4 |
| Gemini-2.0 Flash | 6.5 | 19.7 | 65.8 | |
| Qwen 2.5 72B | 5.6 | 20.5 | 62.0 | |
| LRM + Grounding | Gemini-2.0 Flash Thinking | 2.8 | 14.1 | 59.4 |
| OpenAI o1 | 5.6 | 17.8 | 67.9 | |
| Claude 3.7 Sonnet | 12.1 | 22.8 | 65.4 | |
| Finetuned | UGround-F | 3.6 | 22.8 | 65.7 |
| GPT-4o + UGround-F | 10.3 | 25.0 | 71.0 | |
| Claude 3.7 Sonnet + UGround-F | 11.2 | 25.4 | 68.0 |
@@
Main results are summarized in Table 2. All models fell short in following real-world user instructions. The highest task success rate was only 12.1%, and the highest average progress was only 25.0%. This indicates a big performance gap compared to human annotators, who rarely made errors. Specifically, grounding methods alone failed to follow most instructions. Paired with the best-performing grounding model (UGround-V1), instructions generated by VLMs and LRMs significantly improved the performance. LRMs performed marginally better than most VLMs. When considering all three metrics, GPT-4o, o1, and Claude 3.7 Sonnet have the strongest performance. Finetuning UGround-V1 on real user data significantly improved its performance. However, the benefit is less significant when paired with VLMs and LRMs, since most challenges come from reasoning and planning.
Additionally, we evaluated the best-performing VLM (GPT-4o) + UGround-V1 with varying history context lengths, from no history to full interaction history with the same user, which can be up to 305 steps. An ideal assistant should be able to leverage different kinds of historical context based on different instructions, ranging from no history to multi-task history context (e.g., for routine learning). As shown in Figure 4, increasing context length also does not necessarily lead to better performance. GPT-4o + UGround-V1 achieved the highest task success rate with a context length of 10, and increasing the context length further led to poorer performance. Additionally, providing all past instructions and actions of a user as context is not only expensive, since the context may have hundreds of steps, but also does not increase performance, indicating that the context is not being effectively used. It also suggests a lack of effective routine learning ability.
All baseline experiments used the ground truth transcripts of user speech instructions to avoid introducing errors caused by speech recognition. We provide the results with the speech recognition results as input in Appendix A.1.
5 Discussion
@@
Can grounding models understand real-world user instructions? There remains a significant gap in the performance of current direct grounding methods. All methods show a task success rate of 0, along with low progress and step accuracy. Figure 5 illustrates various failure cases encountered when directly using UGround-V1. Unsurprisingly, grounding models fail to interpret instructions requiring reasoning or contextual understanding, due to their limited reasoning capabilities. However, even for context-free instructions involving straightforward spatial reasoning—tasks where grounding methods should excel—they frequently misinterpret spatial layouts or rankings. For instance, they often incorrectly select elements for instructions such as “click the first one.”
@@
How can VLMs & LRMs help? VLMs or LRMs can convert the original user instructions to more direct and explicit descriptions that a grounding model can more easily understand. This is made possible by their reasoning capacities. For instance, in Figure 5A, the grounding model (UGround-V1) on its own fails to select the first non-sponsored link. However, it succeeds after GPT-4o rewrites the instruction to reference the link’s title directly. As illustrated in Figure 5B, grounding models may sometimes still fail due to inherent limitations even when VLMs/LRMs generate clearer instructions. Nonetheless, incorporating VLMs or LRMs significantly improves overall performance, achieving a non-zero task success rate and much higher progress and step accuracy.
What are the limitations of VLMs & LRMs? While VLMs & LRMs can be helpful, the highest task success rate is still 12.1%. Besides errors caused by grounding models (e.g., Figure 5B), we found that VLMs and LRMs still struggled with complex temporal reasoning. In Figure 5C, the user previously instructed the assistant to open the first two search results in new tabs. When asked to “look at the first one we just opened,” Claude 3.7 Sonnet failed to understand which element “the first one” refers to. Instead of the first tab just opened, it incorrectly referred to the first search result. Evaluation with varying context length further reveals the limited capacity of reasoning from long context in VLMs and LRMs. As shown in Figure 4, model performance does not benefit from having access to longer context. For instance, with full context, models should be able to learn user-specific routines demonstrated in earlier tasks. Despite this, the performance became worse in this setting.
Does learning from real-world user data help? Finetuning UGround-V1 significantly improved the average progress and step accuracy, achieving performance comparable to incorporating VLMs or LRMs. When paired with VLMs or LRMs, performance continues to improve, particularly on task success rate, but the increase in progress and step accuracy is limited. These results suggest that the finetuned model can better understand real user instructions. However, finetuning grounding models alone is insufficient. Current VLMs and LRMs still lack the crucial reasoning and planning abilities to robustly perform the sequential instruction following tasks.
@@
Limitations. Evaluating web agent planning methods, particularly model-based methods (e.g., Putta et al. (2024)), requires interactive simulation environments. Since we focus on understanding real-world user instructions on real-world websites, it is challenging to ensure safety and reproducibility beyond the offline setting adopted in our benchmark and similar prior works (Deng et al., 2024; Cheng et al., 2024). We believe that RealWebAssist is complementary to interactive evaluation benchmarks like WebArena, which focus on planning for a single task. We believe that web agents should be evaluated on both types of benchmarks to fully assess their capabilities. Additionally, while the number of instructions is on par with existing benchmarks, they were collected from 10 participants. We intend to increase user diversity in future versions of the benchmark. Lastly, the current setting does not allow dialogue between a user and the AI assistant.
6 Conclusion
@@
In this paper, we present RealWebAssist, the first benchmark designed to evaluate web agents’ ability to provide long-horizon web assistance with real-world users via sequential instruction-following. Our benchmark poses novel challenges for web assistance, including spatial and temporal reasoning, planning, and learning about user-specific routines. We conducted a comprehensive evaluation and analysis on multiple state-of-the-art GUI grounding models, vision-language models, and large reasoning models. The results reveal critical limitations of current models. We have also shown the benefit of finetuning models on real-world user data. Our benchmark, along with the well-annotated user instruction dataset, provides resources and diagnostic tools for further research on real-world web assistance. In future work, we plan to expand our human study to include more participants from various backgrounds, examine web assistance in interactive settings, and incorporate chat between users and web agents.
7 Acknowledgements
@@
This work was supported by a grant from Amazon.
Appendix A Appendix
A.1 Effect of Speech Recognition Errors
@@
All baseline experiments use the ground truth transcripts of user speech instructions as input to ensure that performance is not affected by errors in speech-to-text transcription. However, in real-world settings, instructions are often given via speech. To reflect this, we evaluated the effect of speech recognition on the agent’s performance by using the transcripts generated from a state-of-the-art automatic speech recognition (ASR) model, Whisper Large-V3 (Radford et al., 2023). Additionally, since users may not always be in quiet, controlled environments using a high-quality microphone like in our user experiment setup, we simulated noisy environments by injecting background noise with noise files from the Microsoft Scalable Noisy Speech Dataset (MS-SNSD) dataset (Reddy et al., 2019), following Ying et al. (2024). The noise files include people talking in the background and keyboard typing sounds. As shown in Table 3, using speech recognition resulted in a 1.9% drop in task success rate, and having noisy speech resulted in a further 1.9% drop. In contrast, the word error rate (WER) of the ASR results increased from 1.4% (original speech) to 28.1% (noisy speech), a much larger performance drop compared to the final task performance. This result suggests that reasoning the true meanings of speech instructions by leveraging context can help mitigate errors from ASR.
@@
| Whisper Large-V3 (with Noise) | 6.5 | 20.6 | 63.4 |
A.2 Dataset Construction Details
Video Segmenting. As shown in the video example A.5, the interactive sessions are highly dynamic, and spoken instructions do not always align cleanly with specific screens or timesteps. Automatically segmenting instructions and matching them to corresponding webpages and actions using heuristics would risk significantly degrading data quality. Therefore, we manually segment the live sessions using video editing software to construct the final RealWebAssist dataset.
@@ A.2 Dataset Construction Details
A.3 More Dataset Statistics
A.3.1 Full List of Tasks
@@
A.3.2 Full List of Websites

### v2

@@ Results
| UI-TARS | 2.8 | 13.1 | 53.8 | |
| GTA-1 | 3.7 | 17.7 | 61.5 | |
| GUI-Actor | 5.7 | 14.7 | 61.4 | |
| VLM + Grounding | GPT-4o + GTA-1 | 8.4 | 23.5 | 72.7 |
| Qwen 2.5 72B + GTA-1 | 9.3 | 24.3 | 69.0 | |
| Gemini 2.5 Flash + GTA-1 | 11.2 | 26.9 | 75.4 | |
| LRM + Grounding | o1 + GTA-1 | 7.5 | 17.7 | 68.2 |
| Gemini 2.5 Pro + GTA-1 | 8.4 | 23.5 | 74.5 | |
| o4-mini + GTA-1 | 10.3 | 21.7 | 67.1 | |
| Claude 3.7 Sonnet + GTA-1 | 12.1 | 26.7 | 68.8 | |
| o3 + GTA-1 | 14.0 | 28.7 | 76.7 | |
| Finetuned | GTA-1-F | 3.7 (+0.0) | 19.7 (+2.0) | 64.3 (+2.8) |
| Gemini 2.5 Flash + GTA-1-F | 11.2 (+0.0) | 26.9 (+0.0) | 75.4 (+0.0) | |
| o3 + GTA-1-F | 14.0 (+0.0) | 28.7 (+0.0) | 76.7 (+0.0) |
@@
Main results are summarized in Table 3. All models fell short in following real user instructions. The highest task success rate was only 14.0%, and the highest average progress was only 28.7%, a large gap compared to humans (93.4% task success rate). This difference has a 95% confidence interval of [71.3, 87.5], and is highly significant with p-value 0.0001. Grounding methods by themselves failed to finish most tasks. However, when paired with the best-performing grounding model (GTA-1), instructions generated by VLMs & LRMs significantly improved the performance. LRMs performed marginally better than most VLMs. Across all three metrics, Gemini 2.5 Flash, Gemini 2.5 Pro, and o3 showed the strongest performance. Finetuning GTA-1 on real user data marginally improved its performance, but finetuning offered no benefit when GTA-1 was paired with VLMs and LRMs, since the finetuned model is trained to adapt to real users’ instructions instead of instructions generated by VLM or LRM.
Discussion
@@
Can grounding models understand real-world user instructions? There remains a significant gap in the performance of current direct grounding methods. The best grounding model, GUI-Actor, has a task success rate of only 5.7%. Figure 5 illustrates various failure cases encountered when directly using GTA-1. Unsurprisingly, grounding models fail to interpret instructions requiring reasoning due to their limited reasoning capabilities. However, even for context-free instructions involving straightforward spatial reasoning—tasks where grounding methods should excel—they frequently misinterpret spatial layouts or rankings. For instance, they often incorrectly select elements for instructions such as “click the first one.”
@@
How can VLMs & LRMs help? VLMs or LRMs can convert the original user instructions into more direct and explicit descriptions that a grounding model can more easily understand. This is made possible by their reasoning capacities. For instance, in Figure 5A, the grounding model (GTA-1) on its own fails to select the first tab: it selects the first element instead of the first tab. However, it succeeds after o3 rewrites the instruction to refer to the title. As shown in Figure 5B, grounding models may sometimes still fail due to inherent limitations even when VLMs/LRMs generate clearer instructions. Nonetheless, incorporating VLMs or LRMs significantly improves overall performance.
What are the limitations of VLMs & LRMs? While VLMs and LRMs help, the highest task success rate is still only 14.0%. Beyond errors from grounding models (e.g., Figure 5B), they continue to struggle with complex temporal reasoning. In Figure 5C, the user previously asked to open the first two search results in new tabs. When later instructed to “look at the first one we just opened,” o3 failed to identify which element “the first one” referred to—instead of the first newly opened tab, it pointed to the first search result. We further analyze the error distribution between reasoning errors (the VLM/LRM mistranslates the instruction and refers to the wrong element) and grounding errors (the rewritten instruction is correct, but the grounding model still fails to click the right element). For the best model (o3 + GTA-1), 43.3% of errors are grounding errors and 56.7% are reasoning errors. This suggests that current VLMs and LRMs still lack the reasoning and planning abilities needed to robustly perform sequential instruction-following tasks.
Does learning from real-world user data help? Fine-tuning GTA-1 marginally improved average progress and step accuracy but yielded no additional benefit when paired with VLMs and LRMs. These results show that the fine-tuned model better understands real user instructions, yet it still fails to generalize to instructions generated by VLMs and LRMs. The experiments suggest that finetuning grounding models on a small set of real user instructions provides minimal benefit, and collecting large-scale real user instructions remains a significant challenge.
@@
Limitations. RealWebAssist represents an important first step towards evaluating web agents on long-horizon, real-user tasks. However, it has several limitations. The first is participant scale and diversity. Collecting real-user data is expensive and time-consuming. The number of participants is comparable to prior works that use expert annotators (lu2024weblinx). However, we intend to increase user diversity in future versions of the benchmark. We will also open-source our data collection tools for community expansion of the dataset. Second, like prior benchmarks on real-world websites (deng2024mind2web; cheng2024seeclick), we constrain our evaluation to an offline setting to ensure reproducibility and safety. This is complementary to benchmarks that focus on interactive evaluation in sandbox environments (e.g., WebArena). We believe that web agents should be evaluated on both types of benchmarks to fully assess their capabilities. Lastly, the current setting does not allow dialogue between a user and the AI assistant, which we will explore in future work.
Conclusion
@@
In this paper, we present RealWebAssist, the first benchmark for evaluating web agents’ ability to provide long-horizon web assistance with real-world users via sequential instruction-following. Our benchmark poses novel challenges, including spatial and temporal reasoning, planning, and adapting to user-specific routines. We conducted a comprehensive evaluation and analysis on multiple state-of-the-art GUI grounding models, VLMs, and LRMs, revealing critical limitations of them. We have also shown the limited benefit of finetuning models on real user data. Our benchmark, along with the well-annotated user instruction dataset, provides resources and diagnostic tools for further research on real-world web assistance. In future work, we plan to expand our human study to include more participants from various backgrounds, examine web assistance in interactive settings, and incorporate chat between users and web agents.
Acknowledgements
@@
This work was supported by a research grant from Amazon. We thank Janice Chen for helpful discussions.
More experiment results
Full VLM & LRM + Grounding results
For the best three grounding models, GTA-1 (yang2025gta1), GUI-Actor (wu2025gui) and UI-TARS (qin2025ui), we test their pairing with all the VLMs and LRMs. Table 3 shows the full results. All the evaluation experiments are run on a single A100 GPU for 20 - 40 minutes. Finetuning GTA-1 model takes 4 hours on 4 A100 GPUs.
| VLM + GTA-1 | GPT-4o + GTA-1 | 8.4 | 23.5 | 72.7 |
| Qwen 2.5 72B + GTA-1 | 9.3 | 24.3 | 69.0 | |
| Gemini 2.5 Flash + GTA-1 | 11.2 | 26.9 | 75.4 | |
| LRM + GTA-1 | Claude 3.7 Sonnet + GTA-1 | 12.1 | 26.7 | 68.8 |
| Gemini 2.5 Pro + GTA-1 | 8.4 | 23.5 | 74.5 | |
| o1 + GTA-1 | 7.5 | 21.1 | 73.1 | |
| o3 + GTA-1 | 14.0 | 28.7 | 76.7 | |
| o4-mini + GTA-1 | 10.3 | 21.7 | 67.1 | |
| VLM + GUI-Actor | GPT-4o + GUI-Actor | 6.5 | 18.0 | 67.0 |
| Qwen 2.5 72B + GUI-Actor | 9.3 | 21.4 | 64.9 | |
| Gemini 2.5 Flash + GUI-Actor | 10.3 | 25.6 | 73.1 | |
| LRM + GUI-Actor | Claude 3.7 Sonnet+ GUI-Actor | 7.5 | 18.5 | 63.9 |
| Gemini 2.5 Pro + GUI-Actor | 9.3 | 24.0 | 73.2 | |
| o1 + GUI-Actor | 7.5 | 17.7 | 68.2 | |
| o3 + GUI-Actor | 12.1 | 27.4 | 74.0 | |
| o4-mini + GUI-Actor | 8.4 | 20.0 | 65.1 | |
| VLM + UI-TARS | GPT-4o + UI-TARS | 6.5 | 20.8 | 67.3 |
| Qwen 2.5 72B + UI-TARS | 7.5 | 21.8 | 63.2 | |
| Gemini 2.5 Flash + UI-TARS | 9.3 | 24.1 | 70.2 | |
| LRM + UI-TARS | Claude 3.7 Sonnet + UI-TARS | 9.3 | 17.5 | 61.5 |
| Gemini 2.5 Pro + UI-TARS | 7.5 | 23.4 | 71.6 | |
| o1 + UI-TARS | 6.5 | 18.5 | 66.0 | |
| o3 + UI-TARS | 12.1 | 27.2 | 72.4 | |
| o4-mini + UI-TARS | 7.5 | 19.4 | 62.5 |
Experiment with different context lengths
We evaluated the best-performing VLM (Gemini 2.5 Flash) + GTA-1 with varying history context lengths, from no history to 20 steps. An ideal assistant should be able to leverage different kinds of historical context based on different instructions, ranging from no history to multi-task history context (e.g., for routine learning). As shown in Figure 6, increasing context length also does not necessarily lead to better performance. Gemini 2.5 Flash + GTA-1 achieved the highest task success rate with a context length of 10, and increasing the context length further led to poorer performance. This suggest the limitation of VLM in effectively utilizing historical context for reasoning.
Effect of Speech Recognition Errors
@@
All baseline experiments use the ground truth transcripts of user speech instructions as input to ensure that performance is not affected by errors in speech-to-text transcription. However, in real-world settings, instructions are often given via speech. To reflect this, we evaluated the effect of speech recognition on the agent’s performance by using the transcripts generated from a state-of-the-art automatic speech recognition (ASR) model, Whisper Large-V3 (radford2023robust). Additionally, since users may not always be in quiet, controlled environments using a high-quality microphone like in our user experiment setup, we simulated noisy environments by injecting background noise with noise files from the Microsoft Scalable Noisy Speech Dataset (MS-SNSD) dataset (reddy2019scalable), following (ying2024siftomrobustspokeninstruction). The noise files include people talking in the background and keyboard typing sounds. As shown in Table 4, using speech recognition resulted in a 1.9% drop in task success rate, and having noisy speech resulted in a further 1.9% drop. In contrast, the word error rate (WER) of the ASR results increased from 1.4% (original speech) to 28.1% (noisy speech), a much larger performance drop compared to the final task performance. This result suggests that reasoning the true meanings of speech instructions by leveraging context can help mitigate errors from ASR.
@@
| Whisper Large-V3 (Noise) | 6.5 | 20.6 | 63.4 |
Dataset Construction Details
Video Segmenting. As shown in the video example Video Example, the interactive sessions are highly dynamic, and spoken instructions do not always align cleanly with specific screens or timesteps. Automatically segmenting instructions and matching them to corresponding webpages and actions using heuristics would risk significantly degrading data quality. Therefore, we manually segment the live sessions using video editing software to construct the final RealWebAssist dataset. All participants provided consent to have their speech recorded and included in this dataset.
@@
More Dataset Details
Evaluation detail
User instructions in RealWebAssist require different operations on the webpage, including clicking, scrolling and typing. We believe that action types other than clicking is trivial (for typing actions, the benchmark includes the step of finding the correct place to type instead of the actual typing process), so we only evaluate click‑type actions with annotated bounding boxes are scored; instructions like “scroll” remain in the history but are not counted in our metrics. Of the 1,885 instructions, 1,412 are scored, yielding 1,714 evaluated action steps (one screenshot per step). Tasks average 17.6 evaluated steps.
User behaviors
Figure 8 shows diverse user behaviors in RealWebAssist not present in previous benchmarks. We include a zip file of the live recordings (including audio) from which the examples are taken.
Information seeking As Figure 8A shows, the user is seeking information from different aspects, like images and ratings, before they make the purchase decision.
Comparing different options Figure 8B shows the process of the user viewing two candidates and finally make the decision between them.
Changing minds In Figure 8C, the user is searching for some immersive dining experience. They are checking different restaurants and frequently change their minds when they see more options.
Trial-and-error As Figure 8D shows, the user has several unsuccessful attempts when searching for men’s fashion week. They refer to previous searchs or initiate new ones to look for what they want.
These diverse behaviors increase the complexity of the web assistance: instead of clearly defined-goals, the user themselves are also actively collecting knowledge to make decisions, which requires web assistant to follow the user’s mind and act accordingly.
Full List of Tasks
@@
Full List of Websites